At a dinner party, 1/5 of the guests were vegetarian. 5/8 of those who eat meat, chose chicken. What fractional part of the guests ate chicken for dinner?
Answer by rfer(16322)

4/5 ate meat
4/5*5/8=20/40=1/2 ate chicken